# Active Vision Might Be All You Need: Exploring Active Vision in Bimanual Robotic Manipulation

### v1

@@ VI Conclusion
In this work, we introduced a novel robotic setup featuring a 7-DoF AV arm, which extends the ALOHA 2 system. Through extensive experiments, we demonstrated that AV can significantly improve imitation learning, particularly in tasks that can benefit from proper selection of the camera perspective. Our results suggest that AV may be sufficient to provide the necessary visual feedback for successful task execution across a broad range of tasks, thereby improving the generalizability of robotic platforms and potentially reducing the need for additional, task-specific camera setups. This approach is inspired by how humans dynamically adjust their perspectives using head, neck, and waist movements to optimize their view during manipulation tasks.
To explore the utility of AV, we conducted an in-depth analysis of its role alongside static and conventional moving (eye-in-hand) camera setups. Our findings revealed that adding more cameras does not necessarily enhance performance, and in some cases, may even complicate the system. While AV shows significant potential, it also introduces complexities that warrant further research, particularly in the development of control architectures capable of managing the decoupled nature of visual feedback and control, the expanded action space, and the method’s susceptibility to distribution shift. Additionally, our results raise important questions regarding the data and training requirements for such systems, highlighting the need for continued investigation into how AV can be integrated into robotic platforms.
Moving forward, we aim to contribute to the development of more generalizable, human-like robotic platforms where task-specific visual data is delivered in a targeted, controlled manner, with redundant information filtered at the sensing level rather than at the computational level. The results presented in this paper, along with the introduction of the open-source AV-ALOHA hardware and software, represent important first steps toward achieving this goal. We hope that this work inspires further research and development in this direction, ultimately leading to more efficient and adaptable robotic systems.
@@ Acknowledgement
We thank Mohnish Gopi for assisting in the development of the VR teleoperation app. We also extend our thanks to Mahdi Naddaf for useful discussions and to Soumyajit Ganguly for helping design 3D prints for the tasks.

### v2

@@ VI Conclusion
In this work, we introduced a novel robotic setup featuring a 7-DoF AV arm. Through extensive experiments, we demonstrated that AV can significantly improve imitation learning, particularly in tasks that can benefit from proper selection of the camera perspective. Our results suggest that AV may be sufficient to provide the necessary visual feedback for successful task execution across a broad range of tasks, thereby improving the generalizability of robotic platforms and potentially reducing the need for additional, task-specific camera setups. This approach is inspired by how humans dynamically adjust their perspectives using head, neck, and waist movements to optimize their view during manipulation tasks.
To explore the utility of AV, we conducted an in-depth analysis of its role alongside static and conventional moving (eye-in-hand) camera setups. While AV shows significant potential, it also introduces complexities that warrant further research, particularly in the development of control architectures capable of managing the decoupled nature of visual feedback and control, the expanded action space, and the method’s susceptibility to distribution shift. Additionally, our results raise important questions regarding the data and training requirements for such systems, highlighting the need for continued investigation into how AV can be integrated into robotic platforms.
Moving forward, we aim to contribute to the development of more generalizable, human-like robotic platforms where task-specific visual data is delivered in a targeted, controlled manner, with redundant information filtered at the sensing level rather than at the computational level. The results presented in this paper, along with open-sourcing AV-ALOHA hardware and software, represent important first steps toward achieving this goal. We hope that this work inspires further research in this direction, ultimately leading to more efficient and adaptable robotic systems.
@@ Acknowledgement
We thank Mohnish Gopi for assisting with the VR app and Soumyajit Ganguly for helping on the design tasks.